OpenStudy (anonymous):

how do you simplify 11 1/12 - 2 2/3

OpenStudy (welshfella):

convert the second fraction to 12ths

OpenStudy (welshfella):

2 2/3 = 2 8/12

OpenStudy (welshfella):

you can also convert to improper fractions so we have 133/12 - 32/12 = 101/12 now e can convert that back to a mixed number

OpenStudy (welshfella):

any questions?

OpenStudy (welshfella):

to convert 2/3 to 12ths we multiply numerator and denominator by 4 2* 4 = 8 - = - 3* 4 = 12